OpenStudy (anonymous):

a ball of mass 0.2kg moving with a velocity of 10 m/s hits a bat normally & returns with the velocity of 20 m/s.If the duration of impact is 0.01 sec find average force exerted by the bat

OpenStudy (anonymous):

600 N

OpenStudy (anonymous):

P'-P=F(T'-T) 0.2kg*30m/s=F*0.01s .solving for F u have F=600 N

OpenStudy (anonymous):

thx a lot guys.. :)